OpenStudy (anonymous):

Describe the transformation from y=2 squared to y=2^(x-8).

OpenStudy (anonymous):

\(y=2^2\) is a number, namely \(y=4\) is there a typo here?

OpenStudy (anonymous):

Yeah, I didn't know how to type what you just put down so I had to write it out the long way.

OpenStudy (anonymous):

is it \[y=2^x\] to \[y=2^{x-8}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

from \(y=f(x)\) to \(y=f(x-8)\) the shift is RIGHT eight units

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw